# Efficiently preparing chiral states via fermionic cooling on bosonic quantum hardware

### v1

@@ Appendix D Single-particle density matrix simulation of free fermions
During free evolution under the Hamiltonian (8) of the main text within a fixed sector of the gauge field $u$, the single particle density matrix evolves under a unitary matrix $W(t)$ by conjugation, $R(t)=W^{\dagger}(t)R(0)W(t)$. This unitary is given by the time ordered exponentiation of the Hamiltonian in matrix form, and can be found numerically by Trotterization or using an ordinary differential equation solver. In Sec. III.2 we use second order Trotterization, making $N_{t}$ steps per cooling cycle. Importantly, the unitary depends on the values of the gauge field which are looked up for its computation.
@@
Depolarizing noise is implemented through stochastic application of Pauli errors before each Trotter step of the time evolution and for each spin ($\{\sigma_{\mathbf{J}},\tau_{\mathbf{J}}\}$) with probability of $\zeta/N_{t}$. The Pauli $\tau_{\mathbf{J}}^{\alpha}$ operators are bilinear in the $c$ fermions (6), such that they correspond to evolution under a quadratic fermionic Hamiltonian. Pauli errors $\sigma_{\mathbf{J}}^{\alpha}$ act on the $b$ Majorana fermions, flipping the signs of two of the three gauge fields incident to site $\mathbf{J}$.
Measuring the flux degrees of freedom within this framework simply requires reading off the product of the values of the $\mathbb{Z}_{2}$ gauge fields around the corresponding plaquettes. These values are well defined at every point in time for each trajectory in these simulations, and their fluctuations can be obtained by averaging over multiple trajectories with stochastic noise. Correcting these fluxes in order to return to the flux-free sector, as described in Appendix A, is done by applying a unitary correction given by a product of Pauli $\sigma_{\mathbf{J}}^{\alpha}$ operators.

### v2

@@ IV.3 Single-particle density matrix simulation of free fermions
During free evolution under the Hamiltonian Eq. (8) of the main text within a fixed sector of the gauge field $u$, the single particle density matrix evolves under a unitary matrix $W(t)$ by conjugation, $R(t)=W^{\dagger}(t)R(0)W(t)$. This unitary is given by the time ordered exponentiation of the Hamiltonian in matrix form, and can be found numerically by Trotterization or using an ordinary differential equation solver. In Sec. “Performance with noise” of the main text we use second order Trotterization, making $N_{t}$ steps per cooling cycle. Importantly, the unitary depends on the values of the gauge field which are looked up for its computation.
@@
Depolarizing noise is implemented through stochastic application of Pauli errors before each Trotter step of the time evolution and for each spin ($\{\sigma_{\mathbf{J}},\tau_{\mathbf{J}}\}$) with probability of $\zeta/N_{t}$. The Pauli $\tau_{\mathbf{J}}^{\alpha}$ operators are bilinear in the $c$ fermions [see Eq. (6)], such that they correspond to evolution under a quadratic fermionic Hamiltonian. Pauli errors $\sigma_{\mathbf{J}}^{\alpha}$ act on the $b$ Majorana fermions, flipping the signs of two of the three gauge fields incident to site $\mathbf{J}$.
Measuring the flux degrees of freedom within this framework simply requires reading off the product of the values of the $\mathbb{Z}_{2}$ gauge fields around the corresponding plaquettes. These values are well defined at every point in time for each trajectory in these simulations, and their fluctuations can be obtained by averaging over multiple trajectories with stochastic noise. Correcting these fluxes in order to return to the flux-free sector, as described in Methods subsection “Removing fluxes in 2D”, is done by applying a unitary correction given by a product of Pauli $\sigma_{\mathbf{J}}^{\alpha}$ operators.
Acknowledgements.
G.K. and E.B. were supported by CRC 183 of the Deutsche Forschungsgemeinschaft (subproject A01), a research grant from the Estate of Gerald Alexander, and ISF-MAFAT Quantum Science and Technology Grant no. 2478/24. M.R. acknowledges the Brown Investigator Award, a program of the Brown Science Foundation, the University of Washington College of Arts and Sciences, and the Kenneth K. Young Memorial Professorship for support.
V Author Contributions
G.K., M.S.R., and E.B. designed the research and contributed to writing the paper. G.K. performed the analytical and numerical calculations.
VI Competing interests
The authors declare no competing interests.
VII Data availability
The data that support the findings of this study are available at https://github.com/kishonyWIS/free_fermions.
VIII Code availability
The code used in this work is available at https://github.com/kishonyWIS/free_fermions.